OpenStudy (anonymous):

how do you solve log base 3 of 1/9 = x ? i don't know how to do fractions :/ so imma be asking a lot cos i really don't get it

OpenStudy (anonymous):

exactly the same way we did \[\log_2(\frac{1}{2})=-1\]

OpenStudy (anonymous):

start with equivalent exponential form \[3^x=\frac{1}{9}\]

OpenStudy (anonymous):

then think of what power you would raise 3 to to get the answer of 1/9

OpenStudy (anonymous):

\[3^2=9\] and so \[3^{-2}=\frac{1}{9}\] and so your answer is \[\log_3(\frac{1}{9})=-2\]

OpenStudy (anonymous):

what if it's the opposite way now ? like , log base 9 of 1/3 = x ?